OpenStudy (richyw):

Invert these relations \[\left|S_{n+}\right\rangle=\cos{\frac{\theta}{2}}\left|S_{z+}\right\rangle+\sin{\frac{\theta}{2}}\left|S_{z-}\right\rangle\] \[\left|S_{n-}\right\rangle=\sin{\frac{\theta}{2}}\left|S_{z+}\right\rangle-\cos{\frac{\theta}{2}}\left|S_{z-}\right\rangle\] I need to find \(\left|S_{z+}\right\rangle\) and \(\left|S_{z-}\right\rangle\). I know this should be really simple but it is throwing me off.

OpenStudy (primeralph):

Combine both and/or use the multiple angle formula.

OpenStudy (richyw):

ok so I tried to combine both, but I still get lost. \[ \left|S_{n+}\right\rangle+\left|S_{n-}\right\rangle= \cos{\frac{\theta}{2}}\left|S_{z+}\right\rangle +\sin{\frac{\theta}{2}}\left|S_{z-}\right\rangle +\sin{\frac{\theta}{2}}\left|S_{z+}\right\rangle -\cos{\frac{\theta}{2}}\left|S_{z-}\right\rangle \] \[ \cos{\frac{\theta}{2}}\left|S_{z+}\right\rangle +\sin{\frac{\theta}{2}}\left|S_{z+}\right\rangle =\cos{\frac{\theta}{2}}\left|S_{z-}\right\rangle -\sin{\frac{\theta}{2}}\left|S_{z-}\right\rangle -\left|S_{n+}\right\rangle-\left|S_{n-}\right\rangle \]

OpenStudy (richyw):

\[\left(\cos{\frac{\theta}{2}} +\sin{\frac{\theta}{2}}\right)\left|S_{z+}\right\rangle =\left(\cos{\frac{\theta}{2}} -\sin{\frac{\theta}{2}}\right)\left|S_{z-}\right\rangle -\left|S_{n+}\right\rangle-\left|S_{n-}\right\rangle\]

OpenStudy (richyw):

yeah I am super lost. here. I could divide both sides by \((\cos\frac{\theta}{2}+\sin{\frac{\theta}{2}})\) but it seems that will get me nowhere.

OpenStudy (primeralph):

http://www.mathsrevision.net/double-angle-formulae

OpenStudy (primeralph):

I'm a little too busy to help now. How good is your trig?

OpenStudy (richyw):

basically my solutions he is just flipping the eigenkets. I just would like to see how this is done.

OpenStudy (primeralph):

Is this Quantum or just math?

OpenStudy (richyw):

he does it again here.

OpenStudy (richyw):

it's quantum mechanics, but it's the math that I don't understand. my trig is OK. Like I could follow any math using trig identities. sometimes I have trouble "Seeing" which identities to use. I can derive them all though.

OpenStudy (primeralph):

Well, if it's Quantum, you'll need more information about the kets.

OpenStudy (richyw):

what information is that?

OpenStudy (primeralph):

Hold on, post the full question.

OpenStudy (richyw):

ok. I can figure out the answer without ever "inverting the relations". but in my solutions he did that and I would like to know how he did it. Here is the question

OpenStudy (richyw):

and here is the solution I was given, with the part I am unclear of highlighted. This same method comes up a lot and I do not understand how he can just "switch" the eigenkets like that.

OpenStudy (richyw):

also the "results from question 2" was just writing \(\left|S_{n+}\right\rangle\) and \(\left|S_{n-}\right\rangle\) in the basis of \(\left|S_{z+}\right\rangle\) and \(\left|S_{z-}\right\rangle\) which I posted initially.

OpenStudy (primeralph):

Can I come back to this later? I'm busy now. Also try posting it in the Physics section. And BUMP it now if you can.

OpenStudy (richyw):

ok thanks a lot. I would really like to know this but for now I can just assume it's always gonna be true on my test! thanks for considering it though!

OpenStudy (anonymous):

There's no quantum mechanics here.

OpenStudy (anonymous):

There's no trigonometry needed for this -- multiply the first equation by cos(theta / 2) and the second by sin(theta / 2). Adding them together will give you your expression for z+. Then multiply the first equation by sin(theta / 2) and the second by cos(theta / 2). Subtracting them gives you the expression for z-

OpenStudy (richyw):

ah. clever. thanks!